6.5 The voltage across a 4-F capacitor is shown in Fig. 6.45. Find the current waveform. 2 4 6 78 ms)
Answer #1

for two points Solution : The formula for straight line equation (x,y) and (21 Y . (y-yo) = YI YI (x-20). . XX, Here, 4= UCHfor oms&t<.8ms The Straight line equation of pornts (mylo)f(86,0) (y-(-(0)) = 0+10 (X=6*103) 1. 87536xio3 19+0 = 5ooo (x=6783- for anset oms PB) = Cotto) (20- 5000+) - Currón) co-sooo) 3 -20110-3 5 - 2010 A for sins te oms ?Ct) = (utro 6) - (- 40+ s